OpenStudy (goformit100):

Fox Mulder is trapped in a maze. To find his way out, he walks 10.0 m, makes a 90.0° right turn, walks 5.00 m, makes another 90.0° right turn, and walks 7.00 m. What is his displacement from his initial position?

OpenStudy (zpupster):

 like this??

OpenStudy (goformit100):

ok ....

OpenStudy (goformit100):

then ?

OpenStudy (zpupster):

I would use Pythagorean theorem and find the hypotenuse a^2 + b^2 = c^2 3^2 + 5^2 = c^2 9+25 = 34 = c^2 c=sqrt34

OpenStudy (goformit100):

ok is it the process originally ?

OpenStudy (zpupster):

that is what I would use

OpenStudy (zpupster):

c is the hypotenuse or the distance(displacement between the 2 points beginning and end

OpenStudy (goformit100):

Thank you I got how to do it

OpenStudy (goformit100):

:)